OpenStudy (anonymous):

what is the second derivative of sec(x+2)

OpenStudy (anonymous):

Do you remember the first derivative of sec(x)? :-)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The derivative of sec(x) = sec(x) tan(x). Now, you have to apply the product rule: derivative of tan(x) = sec^2(x).

OpenStudy (jit4won):

Sec(x+2) x tan(x+2) is the first derivative

OpenStudy (jit4won):

Use product rule

OpenStudy (anonymous):

Because the inside derivative is 1, you don't need to apply the chain rule :-)

OpenStudy (anonymous):

ok is it the product of chain , You giving two different answers

OpenStudy (anonymous):

No, I am talking about sec(x). As you wish to work with x+2, just substitute it. Nothing will be changed, because the inside derivative with respect to x is 1. If you prefer, do the product rule on sec(x+2) tan(x+2).

OpenStudy (anonymous):

ok i will be back with solution to check and see if im on track

OpenStudy (anonymous):

OK then :-)

OpenStudy (anonymous):

I GOT PI^2(SIN(X+2)^2+1/COS(X+2)^3

OpenStudy (anonymous):

Pi? How did you get to that: \[ \frac{d(\sec(x+2)\tan(x+2))}{dx} = \sec(x+2) \frac{d(\tan(x+2))}{dx} + \tan(x+2) \frac{d(\sec(x+2))}{dx}\]

OpenStudy (anonymous):

\( d(\tan(x+2))/dx = \sec^2x \) and \( d(\sec(x+2))/dx = \sec(x+2)\tan(x+2)\)

OpenStudy (anonymous):

idk , the second deriv are a problem the first arent bad

OpenStudy (anonymous):

I mean, sec^2(x+2) for the first sec up there.

OpenStudy (anonymous):

The second derivative is the same thing as taking the first derivative. Shouldn't be any harder :_)

OpenStudy (anonymous):

well calculus period is pointless lol but i till have to take it

OpenStudy (anonymous):

Calculus is one of the culminations of thousands of years of human effort, I would hardly call it pointless. If you need an application, read up on how calculus is used in physics. Hint: it's everywhere.

OpenStudy (anonymous):

well I know it help humans and its history but In real life , not needed

OpenStudy (anonymous):

In real life it's used on a daily basis by professionals in countless fields, all of whom are paid well because they took the time to learn calculus when they were in school.

OpenStudy (anonymous):

Furthermore, learning calculus helps you train your mind to think logically and rigorously, which is a skill that is very hard to gain and is literally priceless.

OpenStudy (anonymous):

The same argument can be applied to playing violin, studying history or economics, doing a sport. By your logic, should we only be concerned about survival? The very fact that we went beyond "real" life is a monument to the human intellect and reasoning, as the violin :-). Abstractions are the corner stone of our age and of what we've achieved, for higher thoughts need a higher language :-)